OpenStudy (anonymous):

@Jaynator495 PREALG4M 3.3.1-5+ Solve the equation. = 3 A. n = 0 B. n = 15 C. n = 8 D. n = 7

OpenStudy (dan815):

whers the equation

OpenStudy (anonymous):

n _ 5

OpenStudy (anonymous):

sry it wont show when I copy and paste

OpenStudy (e.mccormick):

\(\dfrac{n}{5}=3\) OK. So, remember what I showed with multiplying both sides?

OpenStudy (e.mccormick):

All of this has to do with properties of equality: http://www.mathwords.com/e/equation_rules.htm Basically, it all boils down to one concept: What you do to one side of an equality that changes its value, you must do to the other. Otherwise it is no longer an equality. There is one sort of exception. If you multiply one side by something that does not really change it. It rearranges it, but does not change the value.

OpenStudy (anonymous):

sry I was doing something

OpenStudy (e.mccormick):

Well, I'm having a slow day at work, so I get to answer questions here.

OpenStudy (anonymous):

ok

OpenStudy (e.mccormick):

So, what do you think you need to multiply both sides by?

OpenStudy (anonymous):

idk sry im not very good at this

OpenStudy (e.mccormick):

Well, what are you trying to move from one side to the other?

OpenStudy (anonymous):

um im trying to find what number goes where the n should be I think

OpenStudy (anonymous):

oh its 15 bcuz 5*3 is 15

OpenStudy (e.mccormick):

Yes. You multiply both sides by 5 to cancel the \(\dfrac{1}{5}\) on the left. That makes 15 on the right.

OpenStudy (anonymous):

I remember how to do it now I just forgot lol

OpenStudy (e.mccormick):

\(\dfrac{x}{5}\) means the same as \(x\cdot \dfrac{1}{5}\). So you use the multiplicative inverse of \(\dfrac{1}{5}\). That means flip the fractuon upside down! Or \(\dfrac{5}{1}\) wich is just 5.

OpenStudy (anonymous):

ok